I have 5oo bills. some of them are $1.00, some are $100.00, and some are $500.00. When we are added together our result is $50,000.00. How many of each bills do I have.
Answer by rapaljer(4671)

Let x = number of $500 bills
500 - x = number of $100 bills

500x + 100(500 - x ) = 50000
500x + 50000 - 100x = 50000
400x + 50000 = 50000
400x = 0
x= 0

No $500 bills, but all 500 bills are $100 bills, for a total of $50,000.

We should have thought of that before we came up with this fancy equation. Oh, well. . .

R^2 at SCC